How to calculate the number of terms if the first term in an arithmetic series is 2 the last term is 20 and the sum of all terms is 110
Answer by Boreal(15235)

sum=(n/2)(a1+an)
110=(n/2)(2+20)
220=22n
n=10 ANSWER
check
a10=a1+9d
20=2+9d
9d=18
d=2
2,4,6,8,10,12,14,16,18,20